OpenStudy (anonymous):

a galvanized iron sheet 64 cm wide is to be bent to form a gutter by turning up vertical strips along each side.how many centimeters should be turned up at each side in order to give a gutter with the maximum cross-sectional area?

OpenStudy (cwrw238):

2x + y = 64 where x is vertical height of gutter and y = width and cross sectional area = xy are you doing calculus leah?

OpenStudy (cwrw238):

finding derivatives?

OpenStudy (cwrw238):

oh - you're not here - i didn't check lol assuming you have: y = 64 - 2x so lets find area A in terms of x: A = xy = x(64 - 2x) A = 64x - 2x^2 differentiating: dA/dx = 64 - 4x this equals 0 for maxim/minm value of the area A 64 - 4x = 0 so x = 16 this gives maximum area because second derivative is -4 negative so for maxim area the sheet should be turned up 16 cms at each end